Given f(x) = x + 1 and g(x) = x – 1, find (fg)(x).

2x
x^2 – 1
x^2 + 2x – 1
-2

Answer by jim_thompson5910(35256)

(fg)(x) is the same as (f*g)(x) or f(x)*g(x)

This means we are multiplying the functions.

(fg)(x) = (f*g)(x)

(fg)(x) = f(x)*g(x)

(fg)(x) = [ f(x) ] * [ g(x) ]

(fg)(x) = [ x+1 ] * [ x-1 ]

(fg)(x) = x(x-1) + 1(x-1)

(fg)(x) = x^2 - x + x - 1

(fg)(x) = x^2 - 1

So the final answer is choice B) x^2 - 1

Side note: You can use the difference of squares formula as a shortcut.